ACTIVITY: Finding the PD (potential difference). show your solution.

  1. to carry how much charge between two points having potential difference equal to 220V, 1760J of work is done?
  2. the EMF of a cell falls from 3 volts t 2.8 volts when its terminals are joined to n electrical load of 4 Ohms. calculate the internal resistance of the cell.
1
Expert's answer
2021-05-03T07:07:29-0400

1)


"q=\frac{W}{V}=\frac{1760}{220}=8\ C"

2)


"I=\frac{3-2.8}{4}=0.05\ A"